OpenStudy (anonymous):

A circle measures 12 feet in diameter. What’s its area to the nearest foot?

OpenStudy (anonymous):

A of a circle is A=(pi)r^2 A=(pi)(6)^2 \[A\approx113 ft^2\]

myininaya (myininaya):

A=pi*r^2 A=pi*(d/2)^2 A=pi*(12/2)^2 A=pi*(6)^2 A=pi*36 A=36pi